OpenStudy (thedj4jc):

Find the value for the following determinant. [5, -4, 8, 2] = -32 -42 42

OpenStudy (studygurl14):

Do you mean this? \[\left[\begin{matrix}5 & -4 \\ 8 & 2\end{matrix}\right]\]

OpenStudy (thedj4jc):

Yes, it wouldn't let me post it on here.

OpenStudy (studygurl14):

The determinant of a matrix \(\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\) equals \(ad-cb\)

OpenStudy (thedj4jc):

Oh, okay!! Thanks!!!